Ms. Coleman writes today - January 27, 2020, 5 European £ put contracts with the exercise price of CS1.9 = £1.0 and with the expiration date of April 27, 2020. Each £ put size is £31,250. The put premium per £ is CS.02 and the relevant interest rate is 4% per annum

For the following spot exchange rates between CS and £ on April 27, 2020, denoted by ST, calculate profit (loss) of Ms. Coleman of writing or selling 5 £ puts: i ST: CS1.5 per £ ii. Sr: CS 2.1 per £

Also find the break-even ST for Ms. Coleman.
Answer #1
Writing 5 Pound Puts
Each Pound Put Size £        31,250
5 Put size £     156,250
Premium per Pound 0.02 CS
Total Premium received for writing 5 Pound Puts              3,125 (156250*0.02)
Strike Exchange Rate for Put=CS 1.9
Payoff for writing Put: Min.( (ST-1.9),0) CS
ST A B=A*156250 C=B+3125
Spot Rate per Pound (in CS) Payoff per Pound(CS) Total Payoff for 5 Puts(CS) Total Profit/(Loss) in CS
1.5 -0.4          (62,500)            (59,375)
2.1 0 0                3,125
For ST=1.5, Loss          (59,375) CS
For ST=2.1, Profit              3,125 CS
BREAKEVEN ST
For Breakeven ST; Payoff per Pound =-0.02
ST=1.9-0.02= 1.88
BREAKEVEN ST=CS 1.88
ST A B=A*156250 C=B+3125
Spot Rate per Pound (in CS) Payoff per Pound(CS) Total Payoff for 5 Puts(CS) Total Profit/(Loss) in CS
1.88 -0.02            (3,125) 0.00 CS